A boat on a river travels downstream between two points, 60 mi apart, in one hour. The return trip against the current takes 2.5 hours. What is the boat's speed?
___mph
How fast does the current in the river flow?
___ mph


Answer by scott8148(6628)

b + c = 60 mi / 1 hr = 60 mph

b - c = 60 mi / 2.5 hr = 24 mph

adding equations ___ b + c + b - c = 60 + 24 ___ 2b = 84 ___ b = 42

substituting ___ (42) + c = 60 ___ c = 18